Miriam is four years older than ricardo. Ten years ago Miriam was 3 times as old as Ricardo was then. How old are they now?
Answer by stanbon(75887)

Miriam is four years older than ricardo. Ten years ago Miriam was 3 times as old as Ricardo was then. How old are they now?
-------
Equations:
m = r + 4
m-10 = 3(r-10)
-----
Substitute for "m" and solve for "r":
r+4-10 = 3r-30
2r = 24
Ricardo = 12 yrs (age now)
--------------------------
Solve for "m"
m = r + 4
Meriam = 12+4 = 16 yrs (age now)
------------------
cheers,
Stan H.